The information below shows the situation in 2017 and 2018 if the Federal Reserve does not make any change to monetary policy: Year Potential Real GDP Real GDP Price Level 2017 $14 trillion $14 trillion 120 2018 $15 trillion $15.2 trillion 133 a. Compute the economic growth rate and the inflation rate between the two years b. If the Federal Reserve desires to maintain Real GDP in 2018 at the same level of Potential GDP for 2018, what type(s) of policy is the Fed most likely to enact? c. After the Fed enacts the policy change to keep Real GDP in 2018 equal to Potential GDP in 2018, indicate the direction of change (increase/decrease) we would expect in each of the following variables: i. Interest Rates ii. Real GDP iii. Unemployment Rate iv. Inflation Rate
Answer #1

Hi,

Hope you are doing well!

Question:

Answer:

a). Answer:

Data:

Indicators 2017 2018
Potential GDP (USD trillion) 14 15
Real GDP 14 15.2
Price level 120 133

Economic growth rate (2017-18) = [(15.2 - 14)14] *100

= (1.2/14) * 100 = 0.08571*100 = 8.57%

Economic growth rate (2017-18) = 8.57%

Inflation rate = [(133 - 120)120] *100

= (13/120) * 100 = 0.1083 * 100 = 10.83%

Inflation rate = 10.83%

b). Answer:

If the Federal Reserve desires to maintain Real GDP in 2018 at the same level of Potential GDP for 2018, what type(s) of policy is the Fed most likely to enact?

To maintain Real GDP in 2018 at the same level of Potential GDP for 2018 the Fed should follow a follow a expansionary monetary policy (like, increasing discount rate,reserve requirement or QE policy or purchasing government securities through OMO)) that will increase the money supply in the economy that decrease the interest rate. Decreasing interest rate increase the aggregate demand in the economy that boost the growth rate (GDP growth rat). When interest rate decrease its boost the investment level in the economy that boost the employment growth rate and income level.

c). Answer:

After the Fed enacts the policy change to keep Real GDP in 2018 equal to Potential GDP in 2018, indicate the direction of change (increase/decrease) we would expect in each of the following variables: i. Interest Rates ii. Real GDP iii. Unemployment Rate iv. Inflation Rate

To maintain Real GDP in 2018 at the same level of Potential GDP for 2018 the Fed should follow a follow a expansionary monetary policy (like, increasing discount rate,reserve requirement or QE policy or purchasing government securities through OMO)) that will increase the money supply in the economy that decrease the interest rate. Decreasing interest rate increase the aggregate demand in the economy that boost the growth rate (GDP growth rat) and price level. Here interest rate will decrease, real GDP will increase, unemployment rate will decrease and inflation rate increase.

Thank You